What is the yield to maturity on a U.S. Treasury Bill with a face value of $4,500 that is purchased for $4,250?
Answer #1

Consider it as a one year T Bill

YTM or Yield to Maturity = (F-P)/F = (4500-4250)/4500 = 0.055 or 5.55% is the answer to the question